OpenStudy (anonymous):

Amber measured the length of her yard by walking from one end of the yard to the other. She counted 44 steps. Assuming that each of her steps was m in length, Amber said the length of her yard was m. The actual length of her yard is 19 m. What was the percent error in Amber’s measurement? Round to the nearest tenth of a percent, if necessary. A. 13.6% B. 15.8% C. 76% D. 131.6%

OpenStudy (snowsurf):

Your missing where is says m in length.

OpenStudy (anonymous):

its1/2 sorry

OpenStudy (anonymous):

@Jhannybean please help :)

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

im missing a length XD

OpenStudy (anonymous):

Amber measured the length of her yard by walking from one end of the yard to the other. She counted 44 steps. Assuming that each of her steps was 1/2 m in length, Amber said the length of her yard was22 m. The actual length of her yard is 19 m. What was the percent error in Amber’s measurement? Round to the nearest tenth of a percent, if necessary. A. 13.6% B. 15.8% C. 76% D. 131.6%

TheSmartOne (thesmartone):

So each of her steps was 1/2 m. And she took 44 steps. How much is the length of the distance she walked?

OpenStudy (anonymous):

22 m

TheSmartOne (thesmartone):

Correct. And the actual distance was 19 m.

OpenStudy (anonymous):

yeah

TheSmartOne (thesmartone):

And do you know the percent error formula?

TheSmartOne (thesmartone):

@morganlauren11

OpenStudy (anonymous):

|22m-19m| / 19m

OpenStudy (anonymous):

x 100

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (snowsurf):

percent error \[ = \frac{ (\left| Actual - measurment \right|) }{ Actual } 100%\]

TheSmartOne (thesmartone):

Correct @morganlauren11

OpenStudy (anonymous):

i know how how to do it and i get 30%

OpenStudy (snowsurf):

You should put absolute values because you can get negative numbers. In this case it does not but in some cases you can.

OpenStudy (anonymous):

i did that

TheSmartOne (thesmartone):

Ok @morganlauren11 lets do it step by step

TheSmartOne (thesmartone):

What is 22-19=?

OpenStudy (anonymous):

3

TheSmartOne (thesmartone):

Ok now what is